OpenStudy (anonymous):

Can somebody please explain in detail why the roots to x2 (x-squared)+9=0 are 3i and -3i?

OpenStudy (anonymous):

sure

OpenStudy (anonymous):

\[x^2+9=0\\ x^2=-9\\ x=\pm\sqrt{-9}\\ x=\pm\sqrt{9(-1)}\\ x=\pm\sqrt{9}\sqrt{-1}\\ x=\pm3i\]

OpenStudy (anonymous):

Thank you! That helps a lot.

OpenStudy (anonymous):

yw